3. You are a member of Board who chairs an ad committee of reforming taxes on telecommunication services. The local telecom tax es can amount to as much as 25 percent of a consumer’s phone bill. The high rates on telecom services have become quite controversial, due to the fact that the deregulation of the telecom industry has led to a highly competitive market. Your best estimates indicate that, based on current tax rates, the monthly market demand for telecommunication services is given by Q=250-5P and the market supply (including taxes) is Q=4P+110 (both in million). The Board of management is considering tax reform that would dramatically cut tax rates, leading to the supply function under the new tax policy of Q=4.171P+110. How much money would typical consumer save each month as a result of proposed legislation? 


1
Expert's answer
2021-08-16T08:57:46-0400

"Qd=250-5p\\Qs=4P+110"

equilibrium price before tax

"Qd=Qs\\250-5P=4P+110\\4P+5P=250-110\\9P=140\\P=15.55"


Substituting p in demand function we get

"Qd=250-5(15.55)\\Q=172.25units"

supply function under new tax policy

"Qs=4.171+110"

equilibrium price after tax

"Qd=Qs\\250-5P=4.171+110\\250-110=4.171P+5P\\140=9.161P\\P=15.26"


Substituting p in supply function we get

"Qs=4..171(15.26)+110=63.649+110\\Q=173.649 units"

the money per unit saved by typical consumer

"15.55-15.26\\=\$0.29"